OpenStudy (anonymous):

The sum of two numbers is 31. One number is 5 less than 2 times the other number. Find the two numbers.

OpenStudy (anonymous):

Can anyone help me out please?

OpenStudy (anonymous):

(x-5)2y=31 ??

OpenStudy (anonymous):

x+y=31 (x-5)+(2y)=31 I think so, am not sure tho.

OpenStudy (anonymous):

do you know how to solve for y?

OpenStudy (anonymous):

I just tried to solve for Y. I got everything over to the right except the y but didn't look right. I ended up with a fraction

OpenStudy (anonymous):

2y=31-(x-5) 2y=31-x+5 2y=26-x y=(26/2)-x/2 y=13-x/2

OpenStudy (anonymous):

Thats what i got for y.

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

\[\frac{ 13-x }{ 2 }\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

13 minus x/2

OpenStudy (anonymous):

what is X??

OpenStudy (anonymous):

can anyone help me out??

OpenStudy (anonymous):

how about: x-5=31 2y=31

OpenStudy (anonymous):

x=26